OpenStudy (anonymous):

Part a. Write the equation of the line that is parallel to the line y = -2/5x and contains the point (5,-3). Part b. Write the equation of a line that is perpendicular to the line y = 1/4x + 2 and contains the point (11,4).

OpenStudy (anonymous):

y = -2/5x Slope = -2/5 Since the Lines are parallel their slopes are = ie -2/5 using point slope form (5,-3). y-y1=m(x-x1) y + 3 = -2/5 (x - 5) nw rearrange in the form y = mx+c for part b Use the idea that product of their slopes =-1 and use the same idea

OpenStudy (anonymous):

THANK U!!!!!!!!!!!